OpenStudy (anonymous):

(FAN AND MEDAL) help please! A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6, and what does it represent? (4 points)

OpenStudy (anonymous):

@pinkbubbles @ganeshie8????

OpenStudy (ali2x2):

@heretohelpalways Of course it would xD

OpenStudy (ali2x2):

k

OpenStudy (ali2x2):

Are you still there?

OpenStudy (anonymous):

i need help ?

OpenStudy (ali2x2):

Ok

OpenStudy (ali2x2):

well review this and the explanations

OpenStudy (ali2x2):

Part A: Domain will be talking about the values of n. Now when we solve for f(n) = 11.26 we get 7.005... so a reasonable range to plot the graph would be from n = 0 to 8 Part B: The y-intercept is the point where n = 0. But n represents the number of days, i.e. time elapsed. So at n = 0 the point represents the initial reading. Part C: At n = 2, f(2) = 8*(1.05)^2 = 8.82 At n = 6, f(6) = 8*(1.05)^6 = 10.72 The average rate of change is (change in height)/(change in n) The change in height is 10.72 - 8.82 = 1.9 The change in n is 6 - 2 = 4 So the average rate of change is 1.9/4 = 0.475 But what is the average rate of change? It is the average daily (because a single change in n corresponds to a day) growth of the plant (because f(n) represents height) Enjoy :)

OpenStudy (anonymous):

thats the answers?

OpenStudy (ali2x2):

yes

OpenStudy (ali2x2):

You there?

OpenStudy (anonymous):

yes, i just ant to make sure, because this work value 30 pts

OpenStudy (ali2x2):

keystone? what grade are you in?

OpenStudy (ali2x2):

im curious

OpenStudy (anonymous):

now, im 10

OpenStudy (ali2x2):

Ok :P

OpenStudy (anonymous):

u know somebody how can check the answers?

OpenStudy (ali2x2):

well here us this for reviews as well openstudy.com/study#/updates/54047052e4b0f2ed1e143b0f could be useful

OpenStudy (anonymous):

ok thank you!!!

OpenStudy (ali2x2):

No prob :)

OpenStudy (ali2x2):

You can close this question after your done giving me a fan and medal AHEM i mean getting your answers :P